x+y=7 and 2x-y=2 then y=?
2x+y=4 and 3x-2y=13 then x=?

Answer by ewatrrr(24785)

 
Hi
Solution to the following system:
x+y=7
2x-y=2
3x = 9, x = 3 and y = 4
2x+y=4 and ||||multiplying 1st EQ thru by 2 to eliminate the y variable
3x-2y=13
4x+2y=8 and ||||Result of multiplying 1st EQ thru by 2 to eliminate the y variable
3x-2y=13
7x = 21, x = 3 and y = -2 (2x+y = 4 or y = 4-2x)